# Bridging the Linear-Quadratic Gap: A Quantum-Classical Hybrid Approach to Robust Supply Chain Design

### v1

@@ Abstract
The design of supply chain networks in densely populated urban logistics systems faces a timely dilemma: the traditional optimisation approaches are effective to maximise the level of demand perfusion, but they are limited to embracing large expenses in overlapping the facilities and cannibalisation in the market. In the Delhi National Capital Region (NCR), these inefficiencies occur in the form of high operational wastages, which are explained by unnecessary fleet use and overlapping service lines. We resolve this difficulty by redefining the Capacitated Facility Location Problem (CFLP) as a Quadratic Unconstrained Binary Optimisation (QUBO) model and by benchmarking three computational strategies: Classical Greedy heuristics, Exact Branch-and-Bound solvers, and a Quantum-Inspired reverse annealing method. When tested on a high-fidelity digital twin of the Delhi NCR road network of thirty candidate sites, we establish that Classical Greedy algorithms using the theoretical maximum demand of (473 units) lack any theoretical overlap penalty, but incur a prohibitive overlap penalty (5.08). Here, in comparison, the Quantum-Inspired solution only losses 3.2% of demand (450 compared to 465 units relative to the optimal solution), but the solution preserves 21.8% less operational overlap risk (3.26 compared to 4.17), which can be viewed as a 35.8% improvement compared to the Greedy solution. Geospatial analysis shows that it can be attributed to a shift in strategies: This, in contrast to Classical approaches, which focus on locating facilities in the high-density central areas (North/Central Delhi), the quantum-inspired solver autonomously chooses the diversified topology of the North-south network, penetrating into the underserved periphery growth markets in Gurugram and South Delhi. This is a spatially balanced arrangement which is congruent to the polycentric structure of modern time megacities, and displays better stability to volatility in demand. We have shown that quantum-inspired optimisation methods can close the so-called Linear-Quadratic Gap phenomenon, i.e. the systematic inability of greedy methods to capture the actual quadratic interactions between facilities, and offer a way of computing the pathway to operationally robust and risk-optimised supply chain networks in dense urban conditions.
@@ 1.5 Research Objectives and Contributions
1.Mathematical Framework: We develop a novel QUBO Hamiltonian for the CFLP that incorporates geospatial overlap as a quadratic penalty term using an exponential distance-decay function. This formulation is embeddable on quantum annealers and serves as a template for similar logistics optimization problems.
2.Digital Twin Validation: We construct a realistic digital twin of Delhi NCR’s urban logistics network using OpenStreetMap data, incorporating actual road network topology, travel distances via Dijkstra’s shortest-path algorithm, and simulated demand distributions. This provides a operationally grounded testbed that captures real-world constraints like one-way streets and river crossings.
3.Comparative Benchmarking: We provide rigorous comparative analysis across three computational paradigms—Classical Greedy heuristics (industry baseline), Exact Branch-and-Bound solvers (mathematical ground truth), and Quantum-Inspired Reverse Annealing (proposed approach)—demonstrating that quantum-inspired methods achieve near-optimal demand capture while reducing operational risk by 21.8% compared to exact solvers.
4.Strategic Insight Extraction: We demonstrate that the quantum-inspired approach autonomously discovers a geospatially diversified network topology (North-South expansion into Gurugram) that classical methods systematically miss, providing actionable strategic guidance for logistics network design in polycentric megacities.
@@ 2.4 Gap in Existing Research and Our Contribution
Focus on distance minimization without incorporating sophisticated penalty functions for market overlap or risk diversification - Treat facility selection as a coverage problem rather than a optimization challenge - Lack validation against industry-standard heuristics making it difficult to assess practical quantum advantage - Use Euclidean distance metrics that ignore real-world transportation network constraints
@@
1.Adapting Modern Portfolio Theory’s covariance framework to logistics network design, viewing overlap as systematically priced risk.
2.Benchmarking against both naive heuristics (Greedy) and optimal solvers (Branch-and-Bound) to establish a complete performance spectrum.
3.Using actual road network topology and travel distances for operational realism.
4.Providing geospatial visualization that translates mathematical solutions into actionable business strategy.
@@ 3.3 Algorithmic Implementation
  • Constructive Greedy Heuristic (Baseline): Sort all N candidate nodes by demand score $D_{i}$ in descending order; select the top K nodes.
    Computational Complexity: O(N log N) for sorting + O(K) for selection = O(N log N)
    Strategic Profile: This represents the "naive" industry standard employed by logistics companies seeking fast deployment decisions. The algorithm makes no attempt to model pairwise interactions $O_{ij}$ and therefore systematically ignores cannibalization risk.
    Implementation: Python with NumPy array sorting
    Use Case: Establishes the baseline performance achievable with zero sophistication, representing what most companies currently use.
  • Exact Classical Solver (Benchmark): Branch-and-Cut algorithm using Mixed-Integer Linear Programming (MILP) formulation with auxiliary variables $Z_{ij}=x_{i}x_{j}$ to linearize quadratic terms.
    Computational Complexity: Exponential worst-case O(2N), polynomial average-case for small N.
    Strategic Profile: Guarantees the mathematically optimal solution by exhaustively exploring the solution space with intelligent pruning. This serves as our "ground truth" for evaluating approximation quality.
    Implementation: PuLP library (v2.7) with CBC solver backend
    Limitations: Becomes computationally intractable for N > 50 in real-time applications (hours to days of computation time)
  • Quantum-Inspired Reverse Annealing (Proposed): Hybrid quantum-classical approach using Reverse Annealing schedule:
    1.Initialization: Begin with the Greedy solution as a warm start
    2.Reverse Anneal: "Melt" the system state back to s = 0.4 (partial superposition)
    3.Forward Anneal: Re-anneal to s = 1.0 (ground state convergence)
    4.Sampling: Collect $N_{reads}$ = 1000 samples to map the energy landscape
    Computational Complexity: O(N² × $N_{reads}$) for QUBO matrix construction and sampling
    Strategic Profile: Explores the local neighborhood around the greedy optimum using simulated quantum tunneling, seeking "balanced" configurations that the greedy algorithm cannot reach due to its commitment to early decisions.
    Implementation: D-Wave Ocean SDK (dimod v0.12)
    Quantum Hardware Readiness: This implementation is directly portable to D-Wave QPUs (Advantage system) without algorithmic modification. The simulated annealing approach faithfully replicates the quantum annealing schedule and can be viewed as a classical preview of quantum hardware performance.
    Reverse Annealing Rationale: Standard forward annealing starts from a random state, which may require many samples to converge. Reverse annealing leverages the greedy solution’s partial optimality (high demand capture) while allowing the quantum-inspired tunneling mechanism to escape its overlap-blind trap—effectively "refining" rather than "starting from scratch."
@@ 4.1 Quantitative Performance Benchmark
Algorithm Demand Capture Overlap Risk Strategic Profile
@@
Greedy Heuristic 473 5.08 High Reward / High Risk
Exact Solver (PuLP) 465 4.17 Balanced Optimum
Quantum Hybrid 450 3.26 Maximal Efficiency / Robust

### v2

@@ Abstract
The design of supply chain networks in densely populated urban logistics systems faces a timely dilemma: the traditional optimisation approaches are effective to maximise the level of demand perfusion, but they are limited to embracing large expenses in overlapping the facilities and cannibalisation in the market. In the Delhi National Capital Region (NCR), these inefficiencies occur in the form of high operational wastages, which are explained by unnecessary fleet use and overlapping service lines. We resolve this difficulty by redefining the Capacitated Facility Location Problem (CFLP) as a Quadratic Unconstrained Binary Optimisation (QUBO) model and by benchmarking three computational strategies: Classical Greedy heuristics, Exact Branch-and-Bound solvers, and a Quantum-Inspired reverse annealing method. When tested on a high-fidelity digital twin of the Delhi NCR road network of thirty candidate sites, we establish that Classical Greedy algorithms using the theoretical maximum demand of (473 units) lack any theoretical overlap penalty, but incur a prohibitive overlap penalty (5.08). Here, in comparison, the Quantum-Inspired solution only losses 3.2% of demand (450 compared to 465 units relative to the optimal solution), but the solution preserves 21.8% less operational overlap risk (3.26 compared to 4.17), which can be viewed as a 35.8% improvement compared to the Greedy solution. Geospatial analysis shows that it can be attributed to a shift in strategies: This, in contrast to Classical approaches, which focus on locating facilities in the high-density central areas (North/Central Delhi), the quantum-inspired solver autonomously chooses the diversified topology of the North-south network, penetrating into the underserved periphery growth markets. This is a spatially balanced arrangement which is congruent to the polycentric structure of modern time megacities, and displays better stability to volatility in demand. We have shown that quantum-inspired optimisation methods can close the so-called Linear-Quadratic Gap phenomenon, i.e. the systematic inability of greedy methods to capture the actual quadratic interactions between facilities, and offer a way of computing the pathway to operationally robust and risk-optimised supply chain networks in dense urban conditions.
@@ 1.5 Research Objectives and Contributions
1. Mathematical Framework: We develop a novel QUBO Hamiltonian for the CFLP that incorporates geospatial overlap as a quadratic penalty term using an exponential distance-decay function. This formulation is embeddable on quantum annealers and serves as a template for similar logistics optimization problems.
2. Digital Twin Validation: We construct a realistic digital twin of Delhi NCR’s urban logistics network using OpenStreetMap data, incorporating actual road network topology, travel distances via Dijkstra’s shortest-path algorithm, and simulated demand distributions. This provides a operationally grounded testbed that captures real-world constraints like one-way streets and river crossings.
3. Comparative Benchmarking: We provide rigorous comparative analysis across three computational paradigms—Classical Greedy heuristics (industry baseline), Exact Branch-and-Bound solvers (mathematical ground truth), and Quantum-Inspired Reverse Annealing (proposed approach)—demonstrating that quantum-inspired methods achieve near-optimal demand capture while reducing operational risk by 21.8% compared to exact solvers.
4. Strategic Insight Extraction: We demonstrate that the quantum-inspired approach autonomously discovers a geospatially diversified network topology (North-South expansion into Gurugram) that classical methods systematically miss, providing actionable strategic guidance for logistics network design in polycentric megacities.
@@ 2.4 Gap in Existing Research and Our Contribution
Focus on distance minimization without incorporating sophisticated penalty functions for market overlap or risk diversification. Treat facility selection as a coverage problem rather than a optimization challenge. Lack validation against industry-standard heuristics making it difficult to assess practical quantum advantage. Use Euclidean distance metrics that ignore real-world transportation network constraints.
@@
1. Adapting Modern Portfolio Theory’s covariance framework to logistics network design, viewing overlap as systematically priced risk.
2. Benchmarking against both naive heuristics (Greedy) and optimal solvers (Branch-and-Bound) to establish a complete performance spectrum.
3. Using actual road network topology and travel distances for operational realism.
4. Providing geospatial visualization that translates mathematical solutions into actionable business strategy.
@@ 3.3 Algorithmic Implementation
Constructive Greedy Heuristic (Baseline): Sort all $N$ candidate nodes by demand score $D_{i}$ in descending order and select the top $K$ nodes. The computational complexity is $O(N\log N)$ for sorting plus $O(K)$ for selection, resulting in an overall complexity of $O(N\log N)$. Strategically, this represents the “naive” industry standard employed by logistics companies seeking fast deployment decisions. The algorithm makes no attempt to model pairwise interactions $O_{ij}$ and therefore systematically ignores cannibalization risk. The implementation is carried out in Python using NumPy array sorting. This method establishes the baseline performance achievable with zero sophistication, representing what most companies currently use.
Exact Classical Solver (Benchmark): An exact Branch-and-Cut algorithm is used based on a Mixed-Integer Linear Programming (MILP) formulation with auxiliary variables $Z_{ij}=x_{i}x_{j}$ to linearize quadratic terms. The computational complexity is exponential in the worst case, $O(2^{N})$, with polynomial average-case performance for small $N$. Strategically, this approach guarantees the mathematically optimal solution by exhaustively exploring the solution space with intelligent pruning and serves as the “ground truth” for evaluating approximation quality. The implementation uses the PuLP library (v2.7) with the CBC solver backend. However, it becomes computationally intractable for $N>50$ in real-time applications, requiring hours to days of computation time.
Quantum-Inspired Reverse Annealing (Proposed): A hybrid quantum-classical approach is employed using a Reverse Annealing schedule. The process begins with initialization from the greedy solution as a warm start, followed by reverse annealing that partially melts the system state back to $s=0.4$ to induce partial superposition. The system is then forward annealed to $s=1.0$ to drive ground-state convergence, after which $N_{\text{reads}}=1000$ samples are collected to map the energy landscape. The computational complexity is $O(N^{2}\times N_{\text{reads}})$ due to QUBO matrix construction and sampling. This approach explores the local neighborhood around the greedy optimum using simulated quantum tunneling, seeking balanced configurations that the greedy algorithm cannot reach due to its commitment to early decisions. The implementation uses the D-Wave Ocean SDK (dimod v0.12) and is directly portable to D-Wave QPUs (Advantage system) without algorithmic modification. The simulated annealing approach faithfully replicates the quantum annealing schedule and can be viewed as a classical preview of quantum hardware performance. Reverse annealing is motivated by the observation that standard forward annealing starts from a random state and may require many samples to converge; by contrast, reverse annealing leverages the greedy solution’s partial optimality while allowing the tunneling mechanism to escape its overlap-blind trap, effectively refining rather than starting from scratch.
@@ 4.1 Quantitative Performance Benchmark
Algorithm Demand Overlap Strategic
Capture Risk Profile
@@
Greedy 473 5.08 High Reward /
Heuristic High Risk
Exact Solver 465 4.17 Balanced
(PuLP) Optimum
Quantum 450 3.26 Maximal
Hybrid Efficiency